Word problem where the sum equals 9.90 the question is how many are dimes and how many are quarters if there are twice as many dimes as quaters. How do you set up this word problem to arrive at the answers for the numbers of dimes and quarters.
Answer by checkley71(8403)

.10*2x+.25x=9.90
.20x+.25x=990
.45x=9.90
x=9.90/.45
x=22 number of quarters
2*22=44 number of dimes
proof
.10*44+.25*18=990
4.40+4.50=9.90
9.90=9.90